# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v1

@@ 3.2.2. Common Code Patterns in Real-World
After discovering numerous new factors, we then wanted to delve deeper into the prevalent code patterns associated with these issues. Through analysis of Stack Overflow discussions, GitHub issues, and commits, we aim to identify the performance issues commonly faced by developers during development and examine their code patterns. In total, we summarized six common patterns as follows:
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API order, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
@@
Unreleased References Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it will not be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow  [^5] https://stackoverflow.com/questions/56102382, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak.
@@ 3 Private static MySingleton mInstance;
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrades performance. Listing 3 demonstrates a code snippet from Stack Overflow  [^6] https://stackoverflow.com/questions/20942671, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
@@
Large-Scale Data. Large-scale data is also a culprit for performance issues. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow  [^7] https://stackoverflow.com/questions/8820837, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
@@ 5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow  [^8] https://stackoverflow.com/questions/42944376, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw in invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load.
@@
Other Patterns. Besides the previously mentioned patterns, there are several other types of patterns to be aware of, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager in the main thread, etc.
@@
Answers to RQ2.2 We have meticulously identified six common code patterns that contribute to performance issues, which encompass API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
Figure 5. Performance issues and corresponding factors from literature review and real-world data. Colors represent the number of automated tools capable of analyzing each factor (0-5), with darker shades indicating more tools. White indicates no factor-to-performance relationship.

### v2

@@ 2.2.3. Common Code Patterns in Real-World
After identifying key factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues. We summarized six types of common patterns as follows:
@@
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API orders, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
@@ 3 Private static MySingleton mInstance;
Unreleased References. Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it won’t be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak.
@@
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrade performance. Listing 3 demonstrates a code snippet from Stack Overflow, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
@@ 5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
Large-Scale Data. Large-scale data is also a culprit for performance issues. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
@@
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw is invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load.
@@
Other Patterns. Besides the previously mentioned patterns, other types of patterns also warrant attention, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager on the main thread, etc.
@@
Answers to RQ1.3 We identified six prevalent code patterns that commonly contribute to performance issues: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
3. Studies on Performance Issues
In the above Section 2, we systematically examined real-world Android performance issues, their root causes, and common code patterns. However, it remains unclear whether cutting-edge literature targets the same practical issues, whether existing techniques can effectively resolve them, and whether current datasets can comprehensively evaluate the techniques’ effectiveness. To answer these questions, we conducted a systematic literature review (SLR) to assess the current state of Android performance analysis. By comparing the performance issues, techniques, and datasets discussed in research papers with those observed in real-world discussions, we aim to uncover gaps between academic research and real-world challenges.
3.1. Literature Review
To conduct a systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 3 illustrates the working process which comprises 5 steps.
Figure 3. The working process of our systematic literature review.
3.1.1. Keywords Identification
To investigate the status quo of Android performance analysis, we identified a bunch of search keywords to locate relevant publications on this topic. Specifically, the keywords comprise two groups (i.e., G1 AND G2). Each group is made up of several keywords. G1 includes terms related to Android mobile phone-related keywords. G2 includes terms related to performance issues identified in Section 2. Please note that to ensure comprehensive coverage and avoid missing relevant studies, we also included broader terms like “performance” and “issue”. The search string is then constructed by combining them, i.e., g1 AND g2, where g1 and g2 represent the disjunction of keywords from groups G1 and G2, respectively. Please note that these keywords are based on our initial understanding of performance issues and are intended to help us identify “seed” papers. As we review these papers, we could uncover new performance issues, prompting a snowballing process that could lead us to further relevant studies. The * symbol represents any string of characters. For example, *phone* can refer to smartphone, phones, or phone. Similarly, issue* can represent issues or issue. Keywords are as follows:
G1: android, mobile, *phone*
G2: resource, response, energy, memory, launch, freeze, lag, battery, responsiveness, cpu, gpu, storage, latency, performance, issue*
Table 2. CORE A/A* ranked software engineering and mobile venues.
Type Source Research Codes Venues
Journals CORE2020 0803, 0805 TSE, TOSEM, EMSE, JSS, IST, TMC
Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
3.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library. From this initial search, we retrieved a total of 11,442 papers. However, many of these studies were outside the scope of our focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, we exclusively chose those published from conferences and journals rated CORE A*/A [^2] CORE provides conference and journal ratings. For conference: https://portal.core.edu.au/conf-ranks/. For journal: https://portal.core.edu.au/jnl-ranks/ from 2012 to 2024. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the titles of relevant publications. As a result, we successfully identified 441 publications distributed across 23 venues as listed in Table 2. Notably, there is no relevant paper identified in ECSA, FOSSACS, and SEAMS.
3.1.3. Paper Exclusion
To gather relevant papers as possible, we initially considered all returned results from our search. However, not all papers aligned with our focus on performance issues in mobile apps caused by poor coding practices, as discussed in Section 1. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
  1. (1)
  2. (2)
    Papers targeting on non-Android devices (e.g., Windows Phone, IOS devices) were excluded.
  3. (3)
    Papers using Android performance issues as metrics/indicators for comparison were excluded. For instance, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations, and Eprof (Pathak2012WhereIT, ) is a profiler used to monitor energy consumption in apps.
  4. (4)
    Papers targeting Android performance issues but not focused on the mobile app level were excluded. For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope and focused on mobile app issues stemming from poor coding practices by developers.
After applying these exclusion criteria, we retained 49 papers that are closely related to the topic of performance issues analysis at the Android mobile level.
3.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 49 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 66 papers.
3.1.5. Cross Check
Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 66 papers to assess the current state of research and identify gaps in addressing real-world performance issues.
3.2. Results
In this Section, we summarize the current state quo of Android performance analysis by thoroughly reviewing these selected papers. Our aim is to assess whether academic research objectives, tools, and datasets align with the real-world challenges faced by developers and users. Specifically, we evaluate: 1) Targeted performance issues and key factors that researchers focus on (Section 3.2.1), 2) Automatic techniques designed to address Android performance issues (Section 3.2.2), and 3) Datasets designed to evaluate the effectiveness of these automatic techniques (Section 3.2.3).
3.2.1. Result - Targeted Performance Issues
In our literature review, we categorized the performance issues discussed in the collected papers and identified the contributing factors to each issue. This approach allowed us to supplement the performance issues and factors not uncovered in our real-world exploration to obtain a holistic taxonomy. Besides, by comparing these targeted issues and factors in academic research with those observed in real-world discussions, we evaluated how well researchers’ focus aligns with the practical challenges faced by developers and users.
Holistic Taxonomy. Through reviewing 66 research papers, we identified five types of performance issues and 44 contributing factors, marking each in the Lit. column of Table 1 with a checkmark (✓). By incorporating insights from both real-world discussions and research papers, we developed a final taxonomy encompassing seven types of performance issues and 72 factors. By examining these issues and their contributing factors from both real-world data and academic findings, we identified significant gaps between research focus and real-world challenges as follows.
Figure 4. Annual publications on each performance issue.
Gaps between research focus and real-world performance issues. As shown in Fig. 4, we identified five key performance issues from literature: energy consumption, responsiveness, memory consumption, CPU usage, and GPU usage. These performance issues are also prevalent in real-world scenarios, indicating that researchers have made significant efforts to address concerns that are relevant to both developers and users. Despite these efforts, there remains a noticeable imbalance in focus. Among the five issues, energy consumption received the most attention, with 53 of the examined papers (80.30%) focusing on this issue, and this trend is likely to continue. However, despite this focus on energy consumption, Section 2.2 revealed that users are more concerned with responsiveness, while developers prioritize memory consumption. This highlights a discrepancy between the performance issues prioritized by researchers, developers, and users. Furthermore, two performance issues prevalent in real-world scenarios have not been explored in literature: storage consumption and internet data usage. These gaps suggest that researchers should place greater emphasis on the performance issues that are of primary concern to both users and developers in real-world apps.
Gaps in academic research on real-world performance issue factors. By analyzing the cases where both the R-W. and Lit. columns are marked with a checkmark (✓), we found that only 21 out of 49 factors discussed in real-world settings (42.86%) have been studied in literature. It indicates that the majority of performance issues encountered in real-world scenarios remain underexplored in academic research. For instance, memory consumption, which is the top concern for developers, real-world data shows 20 contributing factors, but only 8 (40.00%) have been explored in literature. Similarly, for responsiveness, the top concern for users, 14 real-world factors were identified, but only nine (64.29%) have been studied in literature. Less common issues like CPU usage and GPU usage have seen even less attention, with only 25.00% and 50.00% of their respective real-world root causes being studied in literature. These gaps suggest that many root causes of real-world performance issues still remain unexplored by researchers. Notably, all identified root causes of energy consumption have been addressed in literature, indicating that this area has been extensively studied. This observation aligns with our finding that the majority of research efforts are concentrated on energy consumption.
The presence of research-only factors not seen in real-world discussions. By examining rows where the R-W. column is unmarked (✗) but the Lit. column is marked (✓), we identified 23 factors that have been explored in literature but not observed in the real world. Specifically, we found one unseen factor in memory consumption, four in responsiveness, and 18 in energy consumption. This suggests that literature may uncover factors that did not appear in our real-world exploration, thereby further enriching the taxonomy. Notably, most of these unseen factors pertain to energy consumption, accounting for 90.00% of the factors explored in literature but not observed in the real-world. For example, Li et al. (Li2014MakingWA, ) found that OLED displays consume more energy when displaying light colors than dark colors. Similarly, Cito et al. (Cito2016BatteryawareTI, ) observed that frequent advertisement and analytics (A&A) service requests lead to significant energy waste. On the other hand, there is a substantial number of energy consumption factors that appear only in the literature, yet rarely manifest in real-world applications. This indicates a pronounced disconnection between academic research and the practical challenges faced by practitioners.
Answers to RQ2.1: Through an extensive integration of real-world explorations and thorough literature review, we present a comprehensive taxonomy encapsulating seven primary types of performance issues and 72 contributing factors. A noteworthy finding shows that only 42.86% factors contributing to performance issues have been examined in academic research. A significant misalignment exists in the primary concerns of researchers, developers, and users: researchers concentrate on energy consumption, developers prioritize memory usage, and users are chiefly concerned with responsiveness.
3.2.2. Result - Technique
To address performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the list of available automatic approaches collected from our examined papers. Among the 66 papers examined, eight (12.12%) provided publicly accessible tools designed to detect specific factors contributing to particular performance issues. This small fraction can be attributed to several reasons: some papers lack accessible artifacts, others provide only datasets without the corresponding tool code, and in some cases, the links to the artifacts are invalid. This indicates a significant gap between automatic tools with their practical usage to within our community.
In addition, we found that existing tools target only 17 out of the 72 factors in our taxonomy (Table 1), covering just 23.61%. This indicates that the majority of factors contributing to the seven primary types of performance issues are overlooked by SOTA tools, highlighting a significant gap in addressing these issues comprehensively. Moreover, most automated tools are primarily designed to detect performance issues rather than actively resolve them. Given that these detection tools rely on specific code patterns, we can anticipate that the identified patterns can also be applied to remediate performance issues. We then summarize the approaches as follows:
Tool Approach Technique Stage Performance Issue Target Factor Link
VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Memory Consumption Energy Consumption Activity Resource  (Lu2022DetectingRUTool, )
AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural Code Analysis Detect Memory Consumption Energy Consumption Service  (Song2019ServDroidDSTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Energy Consumption Responsiveness Database Operation Loop  (Lyu2018RemoveRFTool, )
PlumbDroid (Bhatt2020AutomatedRO, ) Static Control-Flow Analysis Patch Generation Detect Fix Responsiveness Resource  (Bhatt2020AutomatedROTool, )
µDroid (Jabbarvand2017DroidAE, ) Dynamic Mutation Testing Power Trace Analysis Detect Energy Consumption Resource Wakelock Loop  (Jabbarvand2017DroidAETool, )
ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Energy Consumption Resource  (Jabbarvand2020AutomatedCOTool, )
Static Analysis. Most existing approaches are based on static analysis to detect or fix performance issues, as shown in Table 3. This is intuitive since static analysis involves reasoning about program behavior without execution, making it lightweight to deploy Specifically, taint analysis, symbolic execution (Lu2022DetectingRU, ), code pattern analysis (Pan2020StaticAC, ; Li2019CharacterizingAD, ; Song2019ServDroidDS, ), and data flow analysis (Lyu2018RemoveRF, ) are commonly employed in rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Lyu2018RemoveRF, ; Bhatt2020AutomatedRO, ).
Dynamic Analysis. There are only two studies that primarily employ dynamic analysis approaches to detect performance bugs. This result aligns with expectations because dynamic analysis involves executing apps in a controlled environment and is unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). As shown in Table 3, both dynamic analysis tools target energy consumption issues, particularly unnecessary energy waste such as resource under-utilization, wakelock misuse, and inefficient loops, which are detectable through abnormal energy consumption monitoring. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
Deep Learning. Despite the growing popularity of deep learning (DL) in software engineering tasks like code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ), surprisingly, only one study applied DL to Android performance analysis. Specifically, Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) observed runtime features dynamically and leveraged DL models to locate energy anomalies. The limited adoption of applying DL to Android performance analysis may due to challenges in collecting large datasets that are sufficient to train DL models on performance issues and their corresponding fixes.
Answers to RQ2.2: We identified eight publicly available tools, primarily using static analysis, dynamic analysis, and deep learning approaches. Existing tools target only 17 of the 72 identified factors to performance issues (23.61%), leaving a substantial number of factors unexamined. Most tools focus on detecting performance issues while neglecting their resolution.
Table 4. A list of datasets used for Android mobile performance analysis, based on our literature review.
Dataset Year # APP Type Performance Issue Target Factor Link
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Energy Consumption Responsiveness Database Operations Loop  (Lyu2018RemoveRFTool, )
Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World Energy Consumption Display Screen  (LinaresVsquez2018MultiObjectiveOOTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World Memoy Consumption Energy Consumption Responsiveness UI Rendering View Inefficient Method  (Habchi2019TheROData, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World Memory Consumption Energy Consumption Resource  (Liu2019DroidLeaksACTool, )
Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World Memory Consumption Energy Consumption UI Rendering Data Type Inefficient Method View Wakelock  (Das2020CharacterizingTETool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Memory Consumption Energy Consumption Resource  (Lu2022DetectingRUTool, )
Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World Energy Consumption Responsiveness UI Rendering  (Janssen2022OnTITool, )
Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World Energy Consumption Responsiveness Advertisements  (Petalotis2024AnESTool, )
3.2.3. Result - Dataset
To collect datasets, several types of papers need to be included: 1) Tool papers that construct datasets to evaluate their designed tools; 2) Empirical studies that collect specific datasets to analyze performance issues, where labeled data can be used to support evaluation. After examining all the public links or websites within the 66 identified studies, we identified 10 (15.15%) unique, available, and open-sourced datasets, as shown in Table 4. Notably, we conducted the examination in October 2024. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. In addition, the number of factors covered by these datasets is quite limited. Out of the 72 factors identified in our taxonomy, only 24 (33.33%) are addressed by existing datasets. This indicates a significant need for constructing more comprehensive datasets. Moreover, there is a considerable imbalance in dataset size. The smallest dataset contains only five apps, while the largest has 324, with an average of 105.7 apps per dataset. This imbalance makes it unrealistic for any single dataset to cover all potential issues comprehensively. Even if some tools are effective at detecting performance issues, they may still face soundness challenges due to false negatives. Therefore, more effort is needed to create datasets that are not only comprehensive but also more representative of the diverse range of performance issues.
Answers to RQ2.3: We identified 10 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 72 factors (33.33%), highlighting a significant gap in tool evaluation. These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 324.